OpenStudy (anonymous):

3/2x - 15 = -12 Solving Equations Containing Fractions

OpenStudy (anonymous):

x-=2

OpenStudy (anonymous):

how

OpenStudy (anonymous):

the latter

OpenStudy (anonymous):

no that is wrong is it \[\frac{3}{2}x-15=-12\] like that?

OpenStudy (anonymous):

add 15 to both sides as a first step you get \[\frac{3}{2}x=-12+15\] or \[\frac{3}{2}x=3\]

OpenStudy (anonymous):

then multiply both sides by the reciprocal of \(\frac{3}{2}\) which is \(\frac{2}{3}\)

OpenStudy (anonymous):

you get \[\frac{2}{3}\times \frac{3}{2}x=3\times \frac{2}{3}\] so \[x=2\]

OpenStudy (anonymous):

if you see a fraction in front of the \(x\), for example if you see \[\frac{3}{4}x=12\] the you always multiply by the reciprocal to get \(x\) by itself in my example if \[\frac{3}{4}x=12\] then \[x=12\times \frac{4}{3}=4\times 4=16\]

OpenStudy (anonymous):

thank you; that was very helpful

OpenStudy (anonymous):

thats what i figured but i wasn't sure